# bibliography/references.bib *[itemize]labelindent=10pt, itemindent=0pt, leftmargin=*

### v1

@@ 3.2. Design choices
As a result of the unordered nature of the recorded hits, we do not use positional encoding for the Transformers. For models involving an encoder-only design, i.e., EncCla and EncReg, padding is used to allow variable length inputs. Note that different events could, for instance, have variable track counts, thus resulting in variable numbers of hits. The following sections will elaborate each model design and approach in detail. A simple depiction is presented in Figure 1, covering Transformer-based model approaches, i.e., EncDec (Figure 1(a)), EncCla (Figure 1(b)) and EncReg (Figure 1(c)).
@@
Figure 1. Intuitive visualisations of the inner-workings of three Transformer model designs, EncDec, EncCla and EncReg, respectively.
@@ 3.2.1. Model 1 - EncDec
The overall idea of this method is similar to the approach used in autoregressive large language models, such as GPT, where the model predicts the next word in a sequence based on the preceding words. In our context, instead of predicting the next word, we predict the next hit from an initial set of hits. The model operates in an autoregressive manner, using previous predictions as inputs for subsequent predictions. This approach enables the model to sequentially build a complete track by iteratively predicting each hit based on the hits predicted so far. In contrast to GPT, we use an encoder-decoder design, which is well-suited for handling the sequential dependencies inherent in track reconstruction. This approach enables the model to sequentially build a complete track by iteratively predicting each hit based on the hits predicted so far.
@@ 3.2.2. Model 2 - EncCla
The main advantage of the model is that it can perform hit-to-track classification in one step, i.e. the inference time of the model itself scales sub-linearly with the number of hits. This may not be the case in practice, as the complexity of the model must increase with the number of hits.
@@ 3.2.3. Model 3 - EncReg
This Transformer is a one-step model, meaning that its time complexity during inference can be rather low, especially when optimisation techniques are implemented. With the help of Flash Attention, the pipeline scales roughly linearly. Its speed makes it also suitable to be used as a refiner network, which regresses track parameters per cluster, i.e., per reconstructed track, and identifies falsely associated hits, increasing the purity of the cluster.
@@ 3.2.4. Model 4 - U-Net
The model processes the entire event as a sparse image, meaning that it labels all hits (coordinates) at once without the need of attention layers or iterative pipelines. Despite their implementation is not straightforward due to the necessity of using sparse convolutions, attention procedures could be considered as a step forward leading to the construction of a modified U-Net architecture as seen previously in some literature (AL_Qurri:2023:IUAM). The usefulness of sparse convolutionals relies on both their conceptual simplicity and computing speed. Studies such as (Abratenko:2021:SSSC) show state-of-the-art performance addressing particle two-dimensional space segmentation scenarios.
@@
Perhaps the most challenging aspect of this U-Net-based approach is the considered preprocessing. In the preprocessing step, when converting the original set of $\{x,y,z\}$ numerical values to the set of $\{i,j,k\}$ to create the sparse tensor of coordinates, a scale factor different from one can be considered. In addition, also during this step, it is possible to carry out an interpolation procedure to add more hits between the original ones. As a consequence, both factors will introduce hyperparameters that need exploration, since a bigger scale factor will demand a higher amount of interpolated data, and vice-versa. However, a fixed scale factor of ×$\times 10$ and also $10$ interpolated hits between each pair of original coordinates have lead to the best results that are presented in this manuscript.
@@ 4.1. Prediction performance
To measure the prediction accuracy of our model design, we consider a modified version of the TrackML score (Kiehn:2019:TrackML), named FitAccuracy score. The TrackML score, just like our FitAccuracy, is a custom metric. In the definition of this the TrackML score, reconstructed tracks with four or more hits are considered, and at least 50% of a reconstructed track’s hits must originate from the same truth particle for that track to be considered for the scoring. The score of a track is the sum of correctly assigned hit weights. REDVID simulations do not generate particles and thus do not consider weights, but tracks. Essentially, we replace the true particle with the true track and consider the weight value of 1 for all hits to arrive at our custom FitAccuracy value. In this fashion, we can have a single comparable scoring for all data sets and model designs. Available scoring for each model is provided in Table 1.
@@ 4.2. Computational performance
REDVID - 10-50 linear tracks EncDec n/a n/a 41 s
@@
REDVID - 10-50 helical tracks EncDec n/a n/a 19 s
@@
REDVID - 50-100 helical tracks EncDec n/a n/a 27 s
@@
TrackML - 10-50 tracks EncDec n/a n/a 16 s
@@
TrackML - 200-500 tracks EncDec n/a n/a -
@@ 5. Discussion - Applying the method
Next, we examine the performance of our different designs. Based on the two collected metric types, i.e., the model prediction accuracy results shown in Figure 3(a) and the computational cost results for CPU and GPU, shown in Figures 3(b) and 3(c), respectively, we can look at how the models perform on different complex tasks. The total execution times provided in Figure 3(d) are given as a range for EncCla, EncReg and U-Net models, while EncDec is represented by wall-clock time collection. This is due to implementation differences. We consider the maximum of CPU-time and GPU-time as the lower bound and the sum of the two (CPU-time and GPU-time) as the upper bound for this range, which is an acceptable estimation.
@@
The first observation is that the Encoder-Decoder model takes the most time due to its autoregressive operation. This model scales linearly with the number of hits, which significantly increases computational time as the task complexity grows. Additionally, its performance declines with higher complexity tasks, specifically those involving more than 100 tracks or the reduced TrackML data. To mitigate excessive training times, this model was no longer used for the TrackML data.
@@ 6. Conclusion and future work
Our main concluding remark relates to the effectiveness of the Transformer architecture for the task of particle tracking. Our experiments show serious potential for single-step processing of event hit data, especially with the encoder-classifier architecture. This architecture is inspired by a single-step translation of hits into track candidates. The low computational effort for the CPU side and the GPU side of event inference and good accuracy, shown in Figure 3, are strong indications for this. The total inference time for the assignment of trace classes to approximately, 10 000 hits per event (level 5) was estimated at 3-6 milliseconds (with or without the use of flash attention). It is not possible to directly compare our inference time with other approaches, as the datasets and computer architectures are not identical. However, we would like to point out again that the Kalman filter pipeline in Ref. required 12 seconds CPU-time, and an optimised Kalman filter algorithm required 1.8 seconds CPU-time per event (Atlas:2019:FTRH). A benchmark for GNNs is (Ju:2021:PGDL) and reports 2.2 seconds wall-clock time (including data transfer to GPUs, pre and post-processing steps) for the full TrackML events using an NVIDIA A100 GPU. This emphasises that the one-step transformer-encoder-classifier architecture is worth a closer look and needs to be further investigated to improve its performance through post-processing and preprocessing steps.
@@
Our single-step models should be supplemented in future research by pre and post-processing steps. Preprocessing involves the formation of clusters from hits, i.e., one could envisage a combination of the encoder-regressor model to form clusters, which are then fed into the classifier model. There are many avenues for ML-based post-processing algorithms. Among these, one could again consider a combined application of the presented methods. For example, as the EncDec model is particularly well suited to the prediction of individual hits, it could be used as a post-processing step, where it could determine for each track candidate whether it is missing hits and add those hits to that track. Another simple addition could be a model that predicts - track candidate by track candidate - which hits are incorrectly assigned. This model could use physically informed loss functions that perform the track parameterisation. It would scale with the number of track candidates (this would still be better than models that scale with the number of hit candidates) and would certainly increase accuracy.
@@ Acknowledgements.
This paper is a result of the collaboration between Radboud University, Nikhef, the University of Twente, SURF, the University of Valencia and the University of Amsterdam. The project has been supported by grants and computing resources from NWO (grant no. 62004546 and grant no. OCENW.GROOT.2019.032), the Dutch national e-infrastructure with the support of the SURF (grant no. EINF-7404), Valencian Community (grant no. CIAICO/2021/184), Valencian Foundation (ValgrAI), Quantum Spain project. The work of R. RdA was supported by PID2020-113644GB-I00 from the Spanish Ministerio de Ciencia e Innovación and by the PROMETEO/2022/69 from the Spanish GVA. The author(s) gratefully acknowledges the computer resources at Artemisa, funded by the European Union ERDF and Comunitat Valenciana as well as the technical support provided by the Instituto de Fisica Corpuscular, IFIC (CSIC-UV).

### v2

@@ 3.3. Design choices
The input sequence to the Transformers is a full event, with the tokens corresponding to hit coordinates (x,y,z), which we do not discretise. As a result of the unordered nature of the recorded hits, we do not use positional encoding for the Transformers. For models involving an encoder-only design, i.e., EncCla and EncReg, padding is used to allow variable length inputs. Note that different events could, for instance, have variable track counts, thus resulting in variable numbers of hits. The following sections will elaborate each model design and approach in detail. A simple depiction is presented in Figure 2, covering Transformer-based model approaches, i.e., EncDec (Figure 2(a)), EncCla (Figure 2(b)) and EncReg (Figure 2(c)).
@@
Figure 2. Intuitive visualisations of the inner-workings of three Transformer model designs, EncDec, EncCla and EncReg, respectively. We showcase in a simplistic manner the hit processing applied by each pipeline. Hits are represented by dots, with gray dots representing no track associations, and dots with the same colour belonging to the same track.
@@ 3.3.1. Model 1 - EncDec
The overall idea of this method is similar to the approach used in autoregressive large language models, such as GPT (Radford:2018:ILUG), where the model predicts the next word in a sequence based on the preceding words. In our context, instead of predicting the next word, we predict the next hit from an initial set of hits. The model operates in an autoregressive manner, using previous predictions as inputs for subsequent predictions. This approach enables the model to sequentially build a complete track by iteratively predicting each hit based on the hits predicted so far. In contrast to GPT, we use an encoder-decoder design, which is well-suited for handling the sequential dependencies inherent in track reconstruction. This approach enables the model to sequentially build a complete track by iteratively predicting each hit based on the hits predicted so far.
@@ 3.3.2. Model 2 - EncCla
The main advantage of the model is that it is a one-step model, assigning all hits to tracks in the full event in a single inference step. This makes its time complexity during inference lower than models that require a full inference step per hit prediction, such as the EncDec model.
@@ 3.3.3. Model 3 - EncReg
Similarly to the EncCla model this is a one-step model, making its time complexity during inference lower than models that require a separate inference step per hit. Its speed makes it also suitable to be used as a refiner network, which regresses track parameters per cluster, i.e., per reconstructed track, and identifies falsely associated hits, increasing the purity of the cluster.
@@ 3.3.4. Model 4 - U-Net
The model processes the entire event as a sparse image, meaning that it labels all hits (coordinates) at once without the need of attention layers or iterative pipelines. Despite their implementation is not straightforward due to the necessity of using sparse convolutions, attention procedures could be considered as a step forward leading to the construction of a modified U-Net architecture as seen previously in some literature (AL_Qurri:2023:IUAM). The usefulness of sparse convolutional operations rely on both conceptual simplicity and computing speed. Studies such as (Abratenko:2021:SSSC) show state-of-the-art performance addressing particle two-dimensional space segmentation scenarios.
@@
Perhaps the most challenging aspect of this U-Net-based approach is the considered preprocessing. In the preprocessing step, when converting the original set of $\{x,y,z\}$ numerical values to the set of $\{i,j,k\}$ to create the sparse tensor of coordinates, a scale factor different from one can be considered. In addition, also during this step, it is possible to carry out an interpolation procedure to add more hits between the original ones. The scaling factor determines the physical spacing between the original points in the three-dimensional data. Conversely, the number of points used for interpolation between consecutive points enhances the amount of information provided to the network. Notably, these parameters were considered exclusively during the training process, leaving the test data unaltered. Different interpolation methods can be employed. In this case, the cubic interpolation method from the SciPy library has been utilised. As a consequence, both factors will introduce hyperparameters that need exploration, since a bigger scale factor will demand a higher amount of interpolated data, and vice-versa. However, a fixed scale factor of ×$\times 10$ and also $10$ interpolated hits between each pair of original coordinates have lead to the best results that are presented in this manuscript.
@@ 4.1. Prediction performance
To measure the prediction accuracy of our model designs, we consider a custom metric, named FitAccuracy score. It is essentially identical to the TrackML score (Kiehn:2019:TrackML), with a small modification that concerns the case of REDVID-generated data sets only. The TrackML score is calculated based on the association of weights to hits in the TrackML data. However, REDVID simulations do not generate particles and thus do not consider weights, but tracks. Essentially, we replace the true particle with the true track and consider the weight value of 1 for all hits to arrive at our custom FitAccuracy value. In this fashion, we can have a single comparable scoring for all data sets and model designs. In the definition of the TrackML score and, consequently, FitAccuracy, reconstructed tracks with four or more hits are considered, and at least 50% of a reconstructed track’s hits must originate from the same truth particle for that track to be considered for the scoring. The score of a track is the sum of correctly assigned hit weights. Available scoring for each model is provided in Table 1.
@@
The models with the best overall performance, EncReg and EncCla, were further analysed to assess their physics performance as functions of two observables in the TrackML 10-50 tracks data set: pseudorapidity $\eta$ [^5] Pseudorapidity is a widely used transformation of $\theta$, the angle relative to the beam axis, and is defined as $\eta\equiv-\ln\left[\tan\left(\frac{\theta}{2}\right)\right]$. and transverse momentum $p_{T}$. To evaluate model performance, two representative metrics were chosen: the FitAccuracy score, as defined in Section 4, and the fake rate, which represents the fraction of predicted tracks that do not correspond to true tracks. The results are presented in Figure 4. Both models exhibit similar trends in their FitAccuracy scores, with EncReg showing marginally better performance at higher $p_{T}$ values. However, the fake rate analysis reveals a more pronounced difference: the encoder classifier (EncCla) performs significantly worse across most slices of phase space, except at high $\eta$.
(a) FitAccuracy versus the transverse momentum $p_{T}$.
(b) FitAccuracy versus pseudorapidity $\eta$.
(c) Fake rate versus transverse momentum $p_{T}$.
(d) Fake rate versus pseudorapidity $\eta$.
Figure 4. Performance metrics as a function of physics observables for the EncCla and EncReg models on the TrackML 10-50 tracks data set, in black and red, respectively. (a) FitAccuracy as a function of the transverse momentum $p_{T}$, (b) FitAccuracy as a function of the pseudorapidity $\eta$, (c) fake rate as a function of the transverse momentum $p_{T}$, and (d) fake rate as a function of the pseudorapidity $\eta$. The uncertainties represent the 68% confidence intervals and were calculated via a bootstrapping procedure. The decreasing trend of EncCla in (d) is due to the increase of bin size with $\eta$, which affects both the number of correctly predicted tracks and the total number of predicted tracks.
@@ 4.2. Computational performance
REDVID 10-50 linear tracks EncDec n/a n/a 41 s
@@
REDVID 10-50 helical tracks EncDec n/a n/a 19 s
@@
REDVID 50-100 helical tracks EncDec n/a n/a 27 s
@@
TrackML 10-50 tracks EncDec n/a n/a 16 s
@@
TrackML 200-500 tracks EncDec n/a n/a -
@@
Table 2 can be used to infer a rough estimate for the scaling of the mean inference GPU-time as a function of the input size. However, a more detailed analysis was conducted for the EncReg-FA model, trained on the 200–500 tracks data set. This model is particularly well-suited for scaling to larger events due to its integration of FlashAttention. The results, presented in Figure 5, demonstrate that while the inference GPU-time does not scale linearly with input size, its growth remains sub-quadratic. Notably, even for events with 8 000 tracks, the mean inference time remains below 200 ms, which is promising for applications to high-luminosity LHC data, where similar track multiplicities are anticipated. It is worth noting, however, that for events with higher track counts, larger models might offer improved performance, albeit at the cost of increased execution time. Consequently, this study should be viewed as a rough demonstration of scaling under the specific constraint of fixed model size.
Figure 5. Mean inference GPU-time for the EncReg-FA model trained on the 200–500 tracks per event data set, plotted as a function of the number of tracks per event.
@@ 5. Discussion - Applying the method
Next, we examine the performance of our different designs. Based on the two collected metric types, i.e., the model prediction accuracy results shown in Figure 6(a) and the computational cost results for CPU and GPU, shown in Figures 6(b) and 6(c), respectively, we can look at how the models perform on different complex tasks. The total execution times provided in Figure 6(d) are given as a range for EncCla, EncReg and U-Net models, while EncDec is represented by wall-clock time collection. This is due to implementation differences. We consider the maximum of CPU-time and GPU-time as the lower bound and the sum of the two (CPU-time and GPU-time) as the upper bound for this range, which is an acceptable estimation. Note that under parallel operation of a given computing platform, the CPU and the GPU activity is not necessarily sequential and the lengthiest of the two would be the least execution duration, hence the consideration of a lower bound.
@@
The first observation is that the Encoder-Decoder model takes the most time due to its autoregressive operation. This model does an inference step per hit, which significantly increases computational time with respect to the other models, which classify all event hits in a single inference step. Additionally, its performance declines with higher complexity tasks, specifically those involving more than 100 tracks or the reduced TrackML data. To mitigate excessive training times, this model was no longer used for the TrackML 200-500 tracks data.
@@ 6. Conclusion and future work
Our main concluding remark relates to the effectiveness of the Transformer architecture for the task of particle tracking. Our experiments show potential for single-step processing of event hit data, especially with the encoder-classifier architecture. This architecture is inspired by a single-step translation of hits into track candidates. The low computational effort for the CPU side and the GPU side of event inference and good accuracy, shown in Figure 6, are strong indications for this. The total inference time for the assignment of track classes to approximately, 10 000 hits per event (level 5) was estimated at 3-6 milliseconds (with or without the use of flash attention). It is not possible to directly compare our inference time with other approaches, as the data sets and computer architectures are not identical. However, we would like to point out again that the Kalman filter pipeline in (Atlas:2019:FTRH) required 12 seconds of CPU-time, and an optimised Kalman filter algorithm required 1.8 seconds of CPU-time per event. A benchmark for GNNs is (Ju:2021:PGDL) and reports 2.2 seconds wall-clock time (including data transfer to GPUs, pre and post-processing steps) for the full TrackML events using an NVIDIA A100 GPU. This emphasises that the one-step transformer-encoder-classifier architecture is worth a closer look and needs to be further investigated to improve its performance through post-processing and preprocessing steps.
@@
Our single-step models should be supplemented in future research by pre and post-processing steps. Pre-processing involves the formation of clusters from hits, i.e., one could envisage a combination of the encoder-regressor model to form clusters, which are then fed into the classifier model. There are many avenues for ML-based post-processing algorithms. Among these, one could again consider a combined application of the presented methods. For example, as the EncDec model is particularly well suited to the prediction of individual hits, it could be used as a post-processing step, where it could determine for each track candidate whether it is missing hits and add those hits to that track. Other additions could be models to detect incorrect hit assignments (per track candidate), to predict track parameters (with track candidate hit set as input) and to utilise physics-informed loss functions for an improved parameterisation. Note that our current solutions are limited to formation of hit clusters matching tracks and do not cover the actual parameterisation of a track function. It would scale with the number of track candidates (this would still be better than models that scale with the number of hit candidates) and would certainly increase accuracy.
@@ Acknowledgements.
This paper is a result of the collaboration between Radboud University, Nikhef, the University of Twente, SURF, the University of Valencia and the University of Amsterdam. The project has been supported by grants and computing resources from NWO (grant no. 62004546 and grant no. OCENW.GROOT.2019.032), the Dutch national e-infrastructure with the support of the SURF (grant no. EINF-7404), Valencian Community (grant no. CIAICO/2021/184), Valencian Foundation (ValgrAI), Quantum Spain project. The work of R. RdA was supported by PID2020-113644GB-I00 from the Spanish Ministerio de Ciencia e Innovación and by the PROMETEO/2022/69 from the Spanish GVA. The author(s) gratefully acknowledges the computer resources at Artemisa, funded by the European Union ERDF and Comunitat Valenciana as well as the technical support provided by the Instituto de Fisica Corpuscular, IFIC (CSIC-UV). The authors would furthermore like to thank Viacheslav Pshenov for his valuable contributions to this work, especially for model performance evaluations.